If 0.600 moles of C6H14 are reacted with 5.7 moles of oxygen and the products are collected in a 4.0 L container at 100.0 .C, what is the total pressure at the end of the reaction?

2 C6H14 (l) + 19 O2 (g) à 12 CO2 (g) + 14 H2O (g)
Answer #1

1. Both are in limiting amount.

Moles of CO2 formed = 6×0.60 = 3.6 mol

Moles of H2O formed = 7×0.60 = 4.2 mol

Total moles formed = 3.6+4.2 =7.8 mol

PV =nRT

P = 7.8×0.0821×373/4.0

= 59.7 atm